# Large deviations for the maximum of a reducible two-type branching Brownian motion

### v2

@@ Abstract.
We consider a two-type reducible branching Brownian motion, defined as a particle system on the real line in which particles of two types move according to independent Brownian motions and create offspring at a constant rate. Particles of type $1$ can give birth to particles of types $1$ and $2$, but particles of type $2$ only give birth to descendants of type $2$. Under some specific conditions, Belloum and Mallein in [4] showed that the maximum position $M_{t}$ of all particles alive at time $t$, suitably centered by a deterministic function $m_{t}$, converge weakly. In this work, we are interested in the decay rate of the following upper large deviation probability, as $t\rightarrow\infty$,
@@
We shall show that the decay rate function exhibits phase transitions depending on certain relations between $\theta$, the variance of the underling Brownian motion and the branching rate.
@@ 1. Introduction and main results
An one-dimensional standard binary branching Brownian motion is a continuous-time particle system on the real line which can be constructed as follows. It starts with one individual located at the origin at time 00 that moves according to a Brownian motion with variance $\sigma^{2}>0$. After an independent exponential time of parameter $\beta$, the initial particle dies and gives birth to 2 children that start at the position their parent occupied at its death. These $2$ children then move according to independent Brownian motions with variance $\sigma^{2}$ and give birth independently to their own children at rate $\beta$. The particle system keeps evolving in this fashion for all time.
In this work, we shall study the two-type reducible branching Brownian motion. This is a particle system on the real line in which particles possess a type in addition with their position. Particles of type $1$ move according to Brownian motions with variance $\sigma^{2}$ and branch at rate $\beta$ into two children of type $1$. Additionally, they give birth to particles of type $2$ at rate $\alpha$. Particles of type $2$ move according to Brownian motions with variance $\sigma_{2}^{2}$ and branch at rate $\beta_{2}$, but cannot give birth to descendants of type $1$. Up to a dilation of time and space, in this work we shall always assume that $\beta_{2}=\sigma_{2}=1$. This model have been studied extensively by many authors; see [7, 14, 15, 17].
With abuse of notation, for all $t\geq 0$, we denote by ${\mathcal{N}}_{t}$ the collection of the individuals alive at time $t$ for both the standard BBM and the two-type reducible BBM. And denote by ${\mathcal{N}}_{t}^{1}$ and ${\mathcal{N}}_{t}^{2}$ the set of particles of type 1 and type 2 at time $t$, respectively. For any $u\in{\mathcal{N}}(t)$ and $s\leq t$, let $X^{u}_{s}$ denote the position at time $s$ of the individual $u$ or its ancestor alive at that time. The maximum of the branching Brownian motion at time $t$ is defined as
@@
To distinguish the two-type reducible BBM and the standard BBM, we shall regard the standard BBM as $(\sigma^{2},\beta)$-BBM to emphasis the variance of the Brownian motion and the branching rate. We also denote by $\mathbb{P}_{\sigma^{2},\beta}$ its law and write $\bar{X}^{u}_{s}$ for the position at time $s$ of the individual $u$ of a $(\sigma^{2},\beta)$-BBM.
For $(\sigma^{2},\beta)$-BBM, the asymptotic behaviour of
@@
Lalley and Sellke [16] characterised $\bar{\omega}$ in terms of a random shift of the Gumbel distribution by introducing the so-called derivative martingale.
Parallel results of (2) for the two-type reducible BBM were obtained in [4]. To state their results, we need to decompose the state space $(\beta,\sigma^{2})\in{\mathbb{R}}_{+}^{2}$ into three regions:
@@
$\lim_{t\to\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}}\theta t\right)=A,$
where if $\beta\leq 1$, or $\beta>1$ and $\sigma^{2}\geq 1$, then
@@
where if $\beta>1$, or $\beta\leq 1,\sigma^{2}\leq 1$, we have
@@
$A=\begin{cases}1-\theta^{2},&\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)},\cr\beta-\frac{\theta^{2}}{\sigma^{2},}&\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4},\cr\frac{2\sqrt{(\beta-1)\left(1-\sigma^{2}\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}},&\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}<\theta^{2}<\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.\end{cases}$
@@
where $C(\theta,\bar{u})$ is a constant depending only on $\theta$ and $\bar{u}$; see [11], [2] and [5]. In this work, we only obtain the exponential decay rates for the large deviation probabilities for the maximum of a two-type inreduciable BBM. It is possible to obtain a form as (3) with the help of the so-called truncated second moment method and the F-KPP equation. Our method may also apply to study the cases of $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}\cap\mathcal{C}_{\text{II}}$ or $(\beta,\sigma^{2})\in\mathcal{C}_{\text{II}}\cap\mathcal{C}_{\text{III}}$ or $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}\cap\mathcal{C}_{\text{III}}$; see [17] for related results.
In the next section, we will present some technical lemmas borrowed from existed literatures. The proofs for Theorems 1.1, 1.2 and 1.3 will be given in Sections 3, 4 and 5, respectively.
@@ 2. Technical lemmas
([13], Lemma 2.1) Let $X$ be a centered Gaussian random variables with variance $\sigma^{2}>0$. Then
@@
Furthermore, for $(\sigma^{2},\beta)$-BBM and $10<x<$ 1, $\mathbb{P}_{\sigma^{2},\beta}-a.s.$,
(6) $\frac{1}{t}\log{\mathcal{N}}_{t}\left(\sqrt{2\beta}x\sigma t)\right)\overset{t\rightarrow\infty}{\longrightarrow}(1-x^{2})\beta,$
where ${\mathcal{N}}_{t}(x)$ denote the number of particles, in the $(\sigma^{2},\beta)$-BBM, alive at time $t$ and positioned in $[tx,\infty)$; see [6].
If $u\in\mathcal{N}_{t}^{2}$, we denote by $T(u)$ the time at which the oldest ancestor of type 2 of $u$ was born. The set $\mathfrak{B}$ of particles of type $2$ that are born from a particle of type $1$, that can be defined as
@@
$\quad=\alpha\int_{0}^{t}e^{\beta s+(t-s)}E\left[g((\sigma B_{u\wedge s}+(B_{u}-B_{u\wedge s}),u\leq t),s)\right]{\rm d}s.$
@@
The first result is just the usual many-to-one lemma. The second one can be found in Proposition 4.1 in [4]. ∎
@@ 3.1. Upper bound
From Lemma 2.3, we have
(7) $\mathbb{P}\left(M_{t}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)$
(8) $=$ $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t},X_{t}^{u}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)$
(9) $\leqslant$ $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1},X_{t}^{u}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)+\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2},X_{t}^{u}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)$
(10) $\leqslant$ $2\max\left[e^{\beta t}\mathbb{P}\left(\sigma B_{t}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right),\alpha\int_{0}^{t}e^{\beta s+(t-s)}\mathbb{P}\left(\sigma B_{s}+B_{t}-B_{s}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)\mathrm{d}s\right]$
(11) =:absent:=:= : $2\max\{I_{1},I_{2}\}.$
Then by (5)
$I_{1}$ $\leq O(1)\cdot\frac{1}{\sqrt{t}}\cdot e^{\beta t}\cdot e^{-\frac{2\beta\theta^{2}t^{2}}{2t}}$
@@
Noticing that
$\operatorname{Var}\left(\sigma B_{s}+B_{t}-B_{s}\right)=\sigma^{2}\cdot s+(t-s),$
we have
(12) $I_{2}$ $\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{2\sigma^{2}\beta\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}\mathrm{d}s$
(13) $=\alpha\int_{0}^{1}e^{f(u)t}\cdot\frac{\sqrt{t}}{\sqrt{2\pi\left((\sigma^{2}-1)u+1\right)}}\mathrm{d}u,$
where
(14) $f_{1}(u)=(\beta-1)u+1-\frac{\sigma^{2}\beta\theta^{2}}{(\left.\sigma^{2}-1\right)u+1}.$
Notice that $f_{1}(1)=\beta(1-\theta^{2})$. Then our task is reduced to know $\max_{0\leq u\leq 1}f(u)$. To end this, let us consider
$f_{1}^{\prime}(u)=\beta-1+\frac{\sigma^{2}\beta\theta^{2}}{\left[\left(\sigma^{2}-1\right)u+1\right]^{2}}\left(\sigma^{2}-1\right).$
We have three cases.
Case I. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\leq 1$, then $u\mapsto f_{1}^{\prime}(u)$ is an increasing function. In this case, by noticing that $\beta\sigma^{2}>1$, one has
$f_{1}^{\prime}(0)=\beta-1+\sigma^{2}\beta(\sigma^{2}-1)=\beta(1-\sigma^{2})+\sigma^{2}\beta\sigma^{2}-1>\beta(1-\sigma^{2})+\sigma^{2}-1\geq 0\cdot$
Thus $f_{1}^{\prime}(u)\geq 0$ for all $u\in[0,1]$. Hence $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and then we have
(15) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\beta\left(1-\theta^{2}\right).$
Case II. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\geq 1,\sigma^{2}\geq 1$, then $f_{1}^{\prime}(u)\geq 0$. Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and hence
(16) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\beta\left(1-\theta^{2}\right).$
Case III. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta>1,\sigma^{2}<1$, then obviously $u\mapsto f_{1}^{\prime}(u)$ is decreasing. Then we have three subcases.
First, if $\theta^{2}<\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}$, which implies $f^{\prime}(1)>0$, then $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and
(17) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\beta\left(1-\theta^{2}\right).$
(18) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq 1-\theta^{2}\beta\sigma^{2}.$
Last, if $\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}$, then $f_{1}^{\prime}(0)\geq 0$ and $f^{\prime}(1)\leq 0$. Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$, where $f_{1}^{\prime}(u_{0})=0$. And one can check that
(19) $u_{0}=\frac{\sqrt{\left|\frac{\sigma^{2}-1}{1-\beta}\right|}\cdot\sigma\theta\cdot\sqrt{\beta}-1}{\left|\sigma^{2}-1\right|},$
and
$f_{1}(u_{0})=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
Therefore,
(20) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
This proves the upper bound. ∎
3.2. Lower bound for Theorem 1.1
Notice that
$\mathbb{P}\left(M_{t}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
$\geqslant$ $\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right),\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)\right\}.$
First, from (3), we have
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
$=\beta\left(1-\theta^{2}\right).$
@@
(21) $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
(22) $=\mathbb{P}\left(\exists u\,\in\mathfrak{B}_{t}:X_{T(u)}^{u}+M_{t-T(u)}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
(23) $=1-\mathbb{P}\left(\forall\,u\in\mathfrak{B}_{t}:X_{T(u)}^{u}+M_{t-T(u)}^{u}<\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
(24) $=1-\mathbb{E}\left[\prod_{u\in\mathfrak{B}_{t}}\varphi_{t}\left(X_{T(u)}^{u},T(u)\right)\right]$
(25) $=1-\mathbb{E}\left[\exp\left(-\sum_{u\in\mathfrak{B}_{t}}-\ln\varphi_{t}\left(X_{T(u)}^{u},T(u)\right)\right)\right],$
@@
$1-\varphi_{t}(x,s)=\bar{u}\left(t-s,\,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x\right)=\mathbb{P}_{1,1}\left(\bar{M}_{t-s}\geq\sqrt{2\sigma^{2}\theta^{2}\beta}\cdot t-x\right).$
@@
and $\left\{M^{u}_{t-T(u)}:u\in\mathfrak{B}_{t}\right\}$ are independent random variables such that $M^{u}_{t-s}$ has the same law as $\bar{M}_{t-s}$ under $\mathbb{P}_{1,1}$. Applying lemma 2.2, we have
@@
(26) $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
@@
(27) $=1-\mathbb{E}\left[\exp\left\{-\alpha\int_{0}^{t}\sum_{u\in{\mathcal{N}}_{s}^{1}}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-X_{s}^{u}\right){\rm d}s\right\}\right]$
@@
(28) $\geq 1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-X_{s}^{u}\right){\rm d}s\right\}\right]$
@@
(29) $\geq 1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right]$

### v4

@@ Abstract.
We consider a two-type reducible branching Brownian motion, defined as a particle system on the real line in which particles of two types move according to independent Brownian motions and create offspring at a constant rate. Particles of type $1$ can give birth to particles of types $1$ and $2$, but particles of type $2$ only give births to descendants of type $2$. Under some specific conditions, Belloum and Mallein in [5] showed that the maximum position $M_{t}$ of all particles alive at time $t$, suitably centered by a deterministic function $m_{t}$, converge weakly. In this work, we are interested in the decay rate of the following upper large deviation probability, as $t\rightarrow\infty$,
@@
We shall show that the decay rate function exhibits phase transitions depending on certain relations between $\theta$, the variance of the underlying Brownian motion and the branching rate.
@@ 1. Introduction and main results
A one-dimensional standard binary branching Brownian motion is a continuous-time particle system on the real line which can be constructed as follows. It starts with one individual located at the origin at time 00 that moves according to a Brownian motion with variance $\sigma^{2}>0$. After an independent exponential time of parameter $\beta$, the initial particle dies and gives birth to 2 children that start at the position their parent occupied at its death. These $2$ children then move according to independent Brownian motions with variance $\sigma^{2}$ and give birth independently to their own children at rate $\beta$. The particle system keeps evolving in this fashion for all time.
In this work, we shall study the two-type reducible branching Brownian motion. This is a particle system on the real line in which particles possess a type in addition with their position. Particles of type $1$ move according to Brownian motions with variance $\sigma^{2}$ and branch at rate $\beta$ into two children of type $1$. Additionally, they give birth to particles of type $2$ at rate $\alpha$. Particles of type $2$ move according to Brownian motions with variance $\sigma_{2}^{2}$ and branch at rate $\beta_{2}$, but cannot give birth to descendants of type $1$. Up to a dilation of time and space, in this work we shall always assume that $\beta_{2}=\sigma_{2}=1$. This model has been studied extensively by many authors; see [7, 14, 15, 17].
With abuse of notation, for all $t\geq 0$, we denote by ${\mathcal{N}}_{t}$ the collection of individuals alive at time $t$ for both the standard BBM and the two-type reducible BBM. And denote by ${\mathcal{N}}_{t}^{1}$ and ${\mathcal{N}}_{t}^{2}$ the set of particles of type 1 and type 2 at time $t$, respectively. For any $u\in{\mathcal{N}}(t)$ and $s\leq t$, let $X^{u}_{s}$ denote the position at time $s$ of the individual $u$ or its ancestor alive at that time. The maximum of the branching Brownian motion at time $t$ is defined as
@@
To distinguish the two-type reducible BBM and the standard BBM, we shall regard the standard BBM as $(\sigma^{2},\beta)$-BBM to emphasize the variance of the Brownian motion and the branching rate. We also denote by $\mathbb{P}_{\sigma^{2},\beta}$ its law and write $\bar{X}^{u}_{s}$ for the position at time $s$ of the individual $u$ of a $(\sigma^{2},\beta)$-BBM. For $(\sigma^{2},\beta)$-BBM, the asymptotic behaviour of
@@
Lalley and Sellke [16] characterized $\bar{\omega}$ in terms of a random shift of the Gumbel distribution by introducing the so-called derivative martingale.
Parallel results of (2) for the two-type reducible BBM were obtained in [5]. To state their results, we need to decompose the state space $(\beta,\sigma^{2})\in{\mathbb{R}}_{+}^{2}$ into three regions (see Figure 1 below for the phase diagram):
@@
We remark here that before the work of Belloum and Mallein [5], Biggins’ work [7] was the first to consider the maximum of multi-type reducible branching random walks. After the work of Belloum and Mallein [5], as $\mathcal{C}_{I},\mathcal{C}_{II},\mathcal{C}_{III}$ does not cover all the possible values for $(\beta,\sigma^{2})\in(0,\infty)^{2}$, Belloum [4] studied the case of $\beta=\sigma^{2}=1$. Later, Ma and Ren [17] studied the other cases and completed the phase diagram.
@@
$\lim_{t\to\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)=A,$
where if $\sigma^{2}\geq 1$, then
@@
where if $\sigma^{2}\leq 1$, we have
@@
$A=\begin{cases}1-\theta^{2},&\theta^{2}\leq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)},\cr\beta-\frac{\theta^{2}}{\sigma^{2},}&\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4},\cr\frac{2\sqrt{(1-\beta)\left(\sigma^{2}-1\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}},&\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}<\theta^{2}<\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.\end{cases}$
@@
where $C(\theta,\bar{u})$ is a constant depending only on $\theta$ and $\bar{u}$; see [11], [2] and [6]. In this work, we only obtain the exponential decay rates for the large deviation probabilities for the maximum of a two-type reducible BBM. It is possible to obtain a form as (3) with the help of the so-called truncated second moment method and the F-KPP equation. Our method may also apply to study the cases of $(\beta,\sigma^{2})\in\partial\mathcal{C}_{\text{I}}\cap\partial\mathcal{C}_{\text{II}}$ or $(\beta,\sigma^{2})\in\partial\mathcal{C}_{\text{II}}\cap\partial\mathcal{C}_{\text{III}}$ or $(\beta,\sigma^{2})\in\partial\mathcal{C}_{\text{I}}\cap\partial\mathcal{C}_{\text{III}}$; see [17] for related results.
In the next section, we will present some technical lemmas borrowed from the existing literature. The proofs for Theorems 1.1, 1.2 and 1.3 will be given in Sections 3, 4 and 5, respectively.
@@ 2. Technical lemmas
([13], Lemma 2.1) Let $X$ be a centered Gaussian random variable with variance $\sigma^{2}>0$. Then
@@
If $u\in\mathcal{N}_{t}^{2}$, we denote by $T(u)$ the time at which the oldest ancestor of type 2 of $u$ was born. The set $\mathfrak{B}$ of particles of type $2$ that are born from a particle of type $1$, which can be defined as
@@
$\quad=\alpha\int_{0}^{t}e^{\beta s+(t-s)}\mathbb{E}\left[g((\sigma B_{u\wedge s}+(B_{u}-B_{u\wedge s}),u\leq t),s)\right]{\rm d}s.$
@@
The first result is just the usual many-to-one lemma. The second can be found in Proposition 4.1 in [5]. ∎
@@
Before going into the details, we would like to give a rough explanation about the general strategy to all proofs. Our method is quite simple and direct. For the upper bound, we just use the first moment argument. Then the task is reduced to understand the birth time $T(u)$ of the particle whose position at time $t$ is larger than $\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t.$ For the lower bound, we may think that oldest ancestor of type $2$ of the maximal particle at time $t$ was born at some time $u_{0}t$ in a position around $x_{0}t$ for some $u_{0}$ and $x_{0}$. So the limit $A$ is consist of two parts: the effort of some type $1$ particles to achieve position $x_{0}t$ at time $u_{0}t$ and the effort that some type $2$ particles to achieve position $\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{0}t$ at time $t-u_{0}t$. The same strategy also apply well to the other two cases.
@@ 3.1. Upper bound
First,
(6) $\mathbb{P}\left(M_{t}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
(7) $=$ $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t},X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
(8) $\leqslant$ $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1},X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)+\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2},X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
[…]
Then by Lemma 2.3 and (5), we have
I $\leq e^{\beta t}\mathbb{P}\left(\sigma B_{t}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
$\leq O(1)\cdot\frac{1}{\sqrt{t}}\cdot e^{\beta t}\cdot e^{-\frac{2\beta\theta^{2}t^{2}}{2t}}$
@@
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
$=\mathbb{P}\left(\exists u\,\in\mathfrak{B}_{t}:X_{T(u)}^{u}+M_{t-T(u)}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
$=1-\mathbb{P}\left(\forall\,u\in\mathfrak{B}_{t}:X_{T(u)}^{u}+M_{t-T(u)}^{u}<\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
$=1-\mathbb{E}\left[\prod_{u\in\mathfrak{B}_{t}}\varphi_{t}\left(X_{T(u)}^{u},T(u)\right)\right]$
$=1-\mathbb{E}\left[\exp\left(-\sum_{u\in\mathfrak{B}_{t}}-\ln\varphi_{t}\left(X_{T(u)}^{u},T(u)\right)\right)\right],$
@@
$1-\varphi_{t}(x,s)=\bar{u}\left(t-s,\,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x\right)=\mathbb{P}_{1,1}\left(\bar{M}_{t-s}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x\right).$
@@
and $\left\{M^{u}_{t-T(u)}:u\in\mathfrak{B}_{t}\right\}$ are independent random variables such that $M^{u}_{t-s}$ has the same law as $\bar{M}_{t-s}$ under $\mathbb{P}_{1,1}$. Applying Lemmas 2.2, we have
@@
(10) $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
@@
(11) $=1-\mathbb{E}\left[\exp\left\{-\alpha\int_{0}^{t}\sum_{u\in{\mathcal{N}}_{s}^{1}}\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-X_{s}^{u}\right){\rm d}s\right\}\right].$
With Lemma 2.3 in hand, applying elementary inequality $1-e^{-x}\leq x$ for $x\geq 0$, we have
@@
(12) II $\leq\mathbb{E}\left[\alpha\int_{0}^{t}\sum_{u\in{\mathcal{N}}_{s}^{1}}\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-X_{s}^{u}\right){\rm d}s\right]$
@@
(13) $=\alpha\int_{0}^{t}{\rm d}s\int_{-\infty}^{+\infty}e^{\beta s}\,\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x\right)\frac{1}{\sqrt{2\pi\sigma^{2}s}}e^{-\frac{x^{2}}{2\sigma^{2}s}}{\rm d}x$